# Evaluating and Mitigating Number Hallucinations in Large Vision-Language Models: A Consistency Perspective

### v1

@@
Abstract
Large vision language models have demonstrated remarkable efficacy in addressing challenges related to both textual and visual content. Nevertheless, these models are susceptible to various hallucinations. In this paper, we focus on a new form of hallucination, specifically termed as number hallucination, which denotes instances where models fail to accurately identify the quantity of objects in an image. We establish a dataset and employ evaluation metrics to assess number hallucination, revealing a pronounced prevalence of this issue across mainstream large vision language models (LVLMs). Additionally, we delve into a thorough analysis of number hallucination, examining inner and outer inconsistency problem from two related perspectives. We assert that this inconsistency is one cause of number hallucination and propose a consistency training method as a means to alleviate such hallucination, which achieves an average improvement of 8% compared with direct finetuning method.
Evaluating and Mitigating Number Hallucinations in Large Vision-Language Models: A Consistency Perspective
Huixuan Zhang and Junzhe Zhang and Xiaojun Wan School of Electronics Engineering and Computer Science, Peking University Wangxuan Institute of Computer Technology, Peking University {zhanghuixuan, junzhezhang}@stu.pku.edu.cn wanxiaojun@pku.edu.cn
1 Introduction
@@
Nevertheless, despite the success of large vision-language models, they still grapple with several challenges, with hallucination being the most severe among them. In this context, hallucination primarily pertains to the discrepancy between the model’s output and the content present in the input image. Among various forms of hallucination, object hallucination, which centers on determining whether an object is present in the image, has been extensively discussed. Numerous methods for evaluation and mitigation have been proposed to tackle this particular challenge(Zhou et al. (2023), Leng et al. (2023), Zhai et al. (2023)). The majority of prior studies employ binary classification questions (answered with ‘yes’ or ‘no’) to assess the model’s performance(Li et al. (2023b), Wang et al. (2023b)).
@@
Input Image Prompt Output
[Uncaptioned image] How many motorcycles are there in this picture? Answer in a single number. 2 (Gold answer is 1.)
[Uncaptioned image] How many persons are there in this picture? Answer in a single number. 3 (Gold answer is 4.)
Table 1: Two examples of number hallucination.
@@
To summarize, our contributions are listed as follows:
  • We introduce a new form of object hallucination, specifically referred to as number hallucination, and present a dataset with 20k pieces of data along with corresponding evaluation results. We point out that all LVLMs we investigated have an average MAE of around $2$ on this dataset, which is disasterous. To the best of our knowledge, our study represents the first comprehensive and quantitative exploration of this particular type of hallucination.
@@
2 Evaluating Number Hallucination in LVLMs
@@ 2.1 Task Definition
2.2 Dataset and Evaluation Metrics
@@
Following this approach, we can construct a dataset with ground truth where each data instance can be denoted as $(i,c,n_{c})$, with $i$ denoting the given image and $c$ denoting the specified object. However, based on our observation, the distribution of answers is not uniform. For instance, in the original dataset, there are about 68% questions with answer $n_{c}=1$. What’s more, the distribution of objects is also non-uniform. To address these biases, we propose double k-uniform sampling. The core of this algorithm is limiting the number of data with object $c$ and answer $n_{c}=n$ not exceeding $k$. The details of our algorithm can be found in Algorithm 1. With this method we construct our evaluation dataset with 20k pieces of data. For details of our algorithm and dataset, please refer to Appendix A.
Evaluation Metrics
@@
In evaluating the counting task, our focus extends to two perspectives. Firstly, we aim for the model to provide correct answers to as many questions as possible. Secondly, we seek minimal errors. To address the first perspective, we treat this problem as a multi-classification task and employ F-scores(macro-F1 and weighted-F1) to evaluate the model’s performance. For the second perspective, we use mean absolute error (MAE) as an evaluation metric. The employment of both metrics provides a more comprehensive and quantitative understanding of the severity of number hallucination. For instance, a model that never accurately counts objects, even with a small MAE, is considered ineffective. Similarly, a model that occasionally counts objects correctly but often makes significant errors should also be deemed problematic. Only when a model that exhibits both high F-scores and low MAE can it be considered free from number hallucinations.
2.3 Evaluation Results
@@
We assess number hallucinations using four of the most advanced large vision-language models (LVLMs) [^2] GPT-4V is not used because this large-scale evaluation is very expensive and time-consuming., which are LLaVA-v1.5-7B(namely LLaVA), LLaVA-v1.5-13B (namely LLaVA-L)(Liu et al., 2023a), InstructBLIP (Dai et al. (2023), equipped with Vicuna-7B) and MiniGPT-v2 (Chen et al. (2023a), equipped with LLaMA-2-chat-7B). The results are shown in Table 2. We also add a random result (randomly select an answer ranging from from $1$ to $10$) as a comparison.
@@
Table 2: Number hallucination evaluation results of different LVLMs.
@@
3 Consistency Analysis from Multiple Perspectives
In addition to the primary counting task mentioned in Section 2.1 (namely primal task), there are also relevant tasks that provide insights into the number hallucination exhibited by a model from different perspectives. Therefore, we can achieve a more comprehensive understanding of number hallucination by analyzing the consistency among the outputs of the related tasks and the primal task. In this context, we mainly analyze two types of consistency, which are inner consistency and outer consistency. Inner consistency represents the consistency of output within the same task format with different settings, while outer consistency represents the consistency between outputs of the primal task and the related task. We take two related tasks and reveal the severe inner and outer inconsistency that both tasks exhibit. This inconsistency is a common problem shared by both tasks, and is a potential cause of number hallucination.
3.1 Binary Classification Questions
@@
Table 3: Evaluation results on binary classification questions. Accuracy is calculated with the ground truth yes or no answers. ‘yes’ ratio is the ratio of yes answers generated by the model.
@@
In this context, inner inconsistency refers to model responding ‘yes’ to even contradictory questions. An example is presented in Table 5 in Appendix B. Despite the evident contradiction between the two questions (different $n$ setting), the model answer ‘yes’ to both. In fact, by comparing prompt I and prompt II, model shows this inner inconsistency at a ratio of 41%, which is very high because a model that randomly answers ‘yes’ or ‘no’ will bear an inner inconsistency of just about 25%. This high ratio reveals a significant challenge that it becomes uncertain what the model’s actual understanding is towards the question. Specifically, if the model answers ‘yes’ to a ground truth, we cannot confidently assert that the model precisely knows the number of objects in the given image. Moreover, even if the model answers ‘no’ to a hallucinated expression, it remains unclear whether the model accurately comprehends the ground truth or not. This inner inconsistency suggests that the model struggles to count the number of objects accurately and exhibits confusion, thereby contributing to number hallucination.
@@ 3.2 Comparison Questions
Table 4: The results of comparison questions. Accuracy is calculated according to ground truth.
@@
Inner inconsistency is similar to what occurs in binary classification questions. In this context, inner inconsistency refers to the model providing different answers to the same question. By comparing within the same prompt format, it is surprising to find that after $c_{1}$ and $c_{2}$ are flipped, the model’s response sometimes changes (see one example in Appendix B in Table 8).
Our quantitative result shows that models generally output inconsistent answers for prompt I at a ratio of 25%, which is non-negligible. Prompt II also experiences this inconsistency even more severely (an example can be found in Appendix B in Table 9). For prompt II, the model responds with inconsistent answers of around 80%, which is disasterous. This inconsistency reveals two facts. Firstly, the order of input and the format of prompt have a notable influence on the model’s output, making it challenging to evaluate the model’s performance with multiple-choice questions. Secondly, this inconsistency also reflects the model’s uncertainty about this question, which is a potential cause of number hallucination.
@@
Table 5: Results of finetuned models.
4 Mitigating Number Hallucination
In this section, we propose a consistency training method to mitigate number hallucination.
4.1 Method
@@ 4.2 Experiment Setup
We utilize the MSCOCO2014 (Lin et al., 2014) training set as the foundational dataset and construct $S$ in the same manner as described in Section 2.2. We refrain from performing k-uniform sampling on the training set to avoid bias and retain as much data as possible. For details about our dataset construction, please refer to Appendix C.
@@
We continue the experiments on LLaVA, InstructBLIP and MiniGPT-v2.444LLaVA and LLaVA-L bear the same architecture, so we only finetune LLaVA for simplicity. For simplicity and efficiency, we freeze the vision encoder and LLM and only finetune the alignment part of each LVLM. Details of our hyper-parameters and models can be found in Appendix C.
4.3 Results and Analysis
The results are presented in Table 5. Some conclusions can be drawn from this result. Firstly, direct finetuning the model is useful for mitigating number hallucinations and provides a fine result.

### v4

# Quantity Matters: Towards Assessing and Mitigating Number Hallucination in Large Vision-Language Models

@@
Abstract.
Large-scale vision-language models have demonstrated impressive skill in handling tasks that involve both areas. Nevertheless, these models frequently experience significant issues with generating inaccurate information, which is hallucination. In this study, we concentrate on a specific type of hallucination—number hallucination, referring to models incorrectly identifying the number of certain objects in pictures. We perform quantitative evaluations regarding number hallucination, showing it to be critical in major open-source large vision-language models. Furthermore, we utilizes two related tasks to conduct an in-depth analysis of number hallucination, revealing the severe inner and outer inconsistency among all tasks. Based on this examination, we devise a training approach aimed at improving consistency to reduce number hallucinations, which leads to an 8% enhancement in performance over direct finetuning methods. Our code and dataset will be released to the community.
large vision-language models, number hallucination, consistency, evaluation, mitigation
copyright: noneccs: Computing methodologies Natural language processingccs: Computing methodologies Computer visionccs: Computing methodologies Computer vision tasks
1. Introduction
@@
Nevertheless, despite the success of large vision-language models, they still grapple with several challenges, with hallucination being the most severe among them. In this context, hallucination primarily pertains to the discrepancy between the model’s output and the content present in the input image. Among various forms of hallucination, object hallucination, which centers on determining whether an object is present in the image, has been extensively discussed. Numerous methods for evaluation and mitigation have been proposed to tackle this particular challenge(Zhou et al. (2023), Leng et al. (2023), Zhai et al. (2023)). The majority of prior studies employ binary classification questions (answered with ‘yes’ or ‘no’) to assess the model’s performance(Li et al. (2023a), Wang et al. (2023a)).
@@
Input Picture Prompt and Output
[Uncaptioned image] Prompt: How many motorcycles are there in this picture? Answer in a single number. Output: 2 Gold Answer: 1
[Uncaptioned image] Prompt: How many persons are there in this picture? Answer in a single number. Output: 3 Gold Answer: 4
Table 1. Two examples of number hallucination.
@@
To summarize, our contributions are listed as follows  [^1] Our code and dataset will be released to the community.:
  • We introduce a new form of object hallucination, specifically referred to as number hallucination, and present a dataset with 20k pieces of data along with corresponding evaluation metrics and results. We point out that all LVLMs we investigated have an average MAE of around $2$ on this dataset, which is disasterous. To the best of our knowledge, our study represents the first comprehensive and quantitative exploration of this particular type of hallucination.
@@
2. Related Works
2.1. Large Vision-Language Models
With the success of large language models (LLMs), there has been a surge in efforts to incorporate additional modalities into LLMs, leveraging their powerful recognition capabilities. BLIP-2 (Li et al., 2023b), for instance, introduces the Q-former to align image input into tokens suitable for utilization by LLMs. LLaVA (Liu et al., 2023b) and LLaVA-v1.5 (Liu et al., 2023a) employ visual instruction tuning and leverage simple linear layers or multi-layer perceptrons to map image output into the text feature space. InstructBLIP (Dai et al., 2023) achieves instruction following abilities by performing instruction tuning on the BLIP-2 backbone. Notably, MiniGPT-4 (Zhu et al., 2023) and MiniGPT-v2 (Chen et al., 2023a) are also powerful LVLMs that exhibit strong performance across various vision-language tasks. Among all open-source and closed-source LVLMs, GPT-4V (OpenAI et al., 2023) retains its status as the most powerful one.
Hallucination is an extensively discussed phenomenon and numerous studies have delved into the examination of hallucinations in large vision-language models (LVLMs). The most extensively discussed form of hallucination is object hallucination, centering around the determination of whether a specific object exists in a given image. In the context of caption-based evaluation, CHAIR (Rohrbach et al., 2018) stands out as a classic metric that compares objects mentioned in the caption with those present in the provided image. POPE (Li et al., 2023a), on the other hand, employs a binary classification approach by directly asking the model whether a particular object exists in the provided image. Zhou et al. (2023) conducted an analysis of factors influencing object hallucination and introduced LURE. Leng et al. (2023) proposed visual contrastive decoding to mitigate object hallucinations, while Zhai et al. (2023) suggested a switch-based method to address this issue. However, these studies predominantly focus on the mere existence of objects, which may not be sufficient for gauging whether the model truly comprehends the image.
Beyond object hallucination, there are also studies addressing more generalized forms of hallucinations. Gunjal et al. (2023) proposed a dataset and reward model to assess and mitigate hallucinations in captions. However, many caption-based hallucination evaluation methods require a lot of training modules, so their performance and stability remain uncertain. RAH (Chen et al., 2023b) and FGHE (Wang et al., 2023a) are two datasets similar to POPE that can be employed for evaluating fine-grained hallucinations. Wang et al. (2023b) and Jing et al. (2023) utilizes LLMs and LVLMs to assist detection of hallucination, but LLMs and LVLMs bear hallucinations themselves, making the evaluation results less reliable. POPE-like evaluation metrics rely on binary classification questions and might not fully capture the model’s ability, making them insufficient for quantitative comparisons.
2.2. Counting Task in VQA
There have been studies specifically addressing the task of counting objects in visual question-answering. Earlier works such as Lempitsky and Zisserman (2010) and Segui et al. (2015) laid some groundwork in this area. Zhang et al. (2018) proposed a counting module. Additionally, Liu et al. (2020) introduced a blockwise counting method to tackle issues like inaccurate generated regression targets and imbalanced samples, which were significant challenges in previous counting methods. However, these methods often require specific designs for the task and modifications to the entire model architecture, which may be insufficient and inefficient in the context of large models nowadays.
For datasets and benchmarks, HowmanyQA (Trott et al., 2017) and TallyQA (Acharya et al., 2019) are two datasets focusing on counting tasks. However, these proposed datasets both bear the problem that both ground truth labels and the queried objects are not uniformly distributed, resulting in probably biased evaluation results. What’s more, these works choose accuracy as one evaluation metric, which is not suitable for non-uniformly distributed labels. Besides, these datasets are directly constructed in the form of VQA, making it hard to be transformed into other complex formats like what we have done in Section 4.2 (because there may be only one counting question regarding one image under most circumstances).
3. Evaluating Number Hallucination in LVLMs
@@ 3.1. Task Definition
3.2. Dataset and Evaluation Metrics
@@
Following this approach, we can construct a dataset with ground truth where each data instance can be denoted as $(i,c,n_{c})$, with $i$ denoting the given image and $c$ denoting the specified object. However, based on our observation, the distribution of answers is not uniform. For instance, in the original dataset, there are about 68% questions with answer $n_{c}=1$. What’s more, the distribution of objects is also non-uniform. To address these biases, we propose double k-uniform sampling. The core of this algorithm is limiting the number of data with object $c$ and answer $n_{c}=n$ not exceeding $k$. We present double k-uniform sampling in algorithm 1. Adjusting the value of $k$ allows us to strike a balance between achieving a more uniform distribution and retaining a larger dataset. In the subsequent discussion, we opt for $k=50$.
Algorithm 1 Double k-Uniform Sampling
1:  Input dataset $S=\{(i,c,n_{c})\}$ and $k$.
2:  Randomly shuffle the dataset $S$.
3:  Initialize $m(c,n)\leftarrow 0,S^{{}^{\prime}}=\phi$
4:  for $j=1,2...,|S|$ do
5:     if $m(c_{j},n_{cj})>k$ then
6:        continue
7:     end if
8:     $S^{{}^{\prime}}\leftarrow S^{{}^{\prime}}\cup\{(i_{j},c_{j},n_{cj})\}$
9:     $m(c_{j},n_{cj})\leftarrow m(c_{j},n_{cj})+1$
10:  end for
11:  Output $S^{{}^{\prime}}$
It is worth mentioning that, when constructing dataset, we only retain data $(i,c,n_{c})$ with $n_{c}>0$. There are two main considerations for this choice. First of all, there are many objects that do not appear in a given image and retaining all these questions is too expensive. If we perform k-uniform sampling similarly, this may result in random selection, which is insufficient for reflecting the model’s performance according to (Li et al., 2023a). More importantly, we conduct an experiment on LLaVA with randomly selected $(i,c,n_{c})$ with $n_{c}=0$ and find out that LLaVA correctly answers 95% of these questions, which meets the its high POPE evaluation result (Liu et al., 2023a). Based on these considerations, we only retain data $(i,c,n_{c})$ with $n_{c}>0$ to focus on the number of an existing object.
Another consideration is that, one bounding box in MSCOCO annotation may not exactly correspond to one existing object. Luckily, we found one solution of utilizing the ‘is_crowd’ field of the annotation. When constructing original dataset $S$, if there exists one annotation $(b_{j},c_{j})$ in image $i$ with ‘is_crowd’ being True, the data point $(i,c_{j},n_{c_{j}})$ is not included in original dataset $S$. This approach ensures the calculation of objects in the original dataset $S$ is correct.
After performing $k$-uniform sampling, we manually sampled and checked some labels in the final $S^{{}^{\prime}}$ to ensure the quality of the final dataset. Finally we obtain a dataset with 20200 pieces of data. The distribution of labels are listed in Figure 1.
Figure 1. The distribution of the answers to the questions.
As can be seen from the Figure 1, though k-uniform sampling is performed, questions with answer $\geq 10$ is still too less. Therefore to avoid randomness, when calculating F-scores, questions with answers $\geq 10$ are considered one category.
Evaluation Metrics and Decoding Strategy
In evaluating the counting task, our focus extends to two perspectives. Firstly, we aim for the model to provide correct answers to as many questions as possible. Secondly, we seek minimal errors.
To address the first perspective, we treat this problem as a multi-classification task and employ F-scores(macro-F1 and weighted-F1) to evaluate the model’s performance. Macro-F1 is more suitable for non-uniformly distributed labels like the ones in our test set because it gives equal weight to both large and small categories. However, since there is less data in some categories, the value of macro-F1 may be more easily affected by the results of these data, which introduces more randomness. To address this problem, weighted-F1 is introduced to offer a more comprehensive understanding of the performance of models. Weighted-F1 gives larger weights to larger categories, which reduces randomness. However, if the model perform much better on larger categories than on smaller ones, weighted-F1 may be cheated. Thus only by combining macro-F1 and weighted-F1 can we fully reveal the capability of models.
@@
For the second perspective, we use mean absolute error (MAE) as an evaluation metric. The employment of both F-scores and MAE provides a more comprehensive and quantitative understanding of the severity of number hallucination. For instance, a model that never accurately counts objects, even with a small MAE, is considered ineffective. Similarly, a model that occasionally counts objects correctly but often makes significant errors should also be deemed problematic. Only when a model that exhibits both high F-scores and low MAE can it be considered free from number hallucinations.
We use greedy decoding to make sure no randomness is introduced to the responses. All following evaluations use the same setting. Since the answers to our questions are relatively short, we believe this setting does not harm performance.
3.3. Evaluation Results
@@
We assess number hallucinations using four of the most advanced large vision-language models (LVLMs), which are LLaVA-v1.5-7B(namely LLaVA), LLaVA-v1.5-13B (namely LLaVA-L)(Liu et al., 2023a), InstructBLIP (Dai et al. (2023), equipped with Vicuna-7B (Chiang et al., 2023)) and MiniGPT-v2 (Chen et al. (2023a), equipped with LLaMA-2-chat-7B (Touvron et al., 2023)). The results are shown in Table 2. We also add a random result (randomly select an answer ranging from from $1$ to $10$) as a comparison. Since testing GPT-4V is very expensive, we perform a quantitative evaluation on a small subset of our test set. Details can be found in Appendix A.
@@
Table 2. Number hallucination evaluation results of different LVLMs.
@@
4. Consistency Analysis from Multiple Perspectives
In addition to the primary counting task discussed in Section 3.1 (namely primal task), there are also relevant tasks that provide insights into number hallucination exhibited by a model. Further, by analyzing the consistency among the outputs of the related tasks and the primal task, we can achieve a more comprehensive understanding of number hallucination. In this context, we mainly analyze two types of consistency, which are inner consistency and outer consistency. Inner consistency represents the consistency of output within the same task format with different settings, while outer consistency represents the consistency between outputs of the primal task and the related task. We take two related tasks and reveal the severe inner and outer inconsistency that both tasks exhibit. This inconsistency is a common problem shared by both tasks, and is a potential cause of number hallucination.
4.1. Binary Classification Questions
@@
Table 3. Evaluation results on binary classification questions. Accuracy is calculated with the ground truth yes or no answers. ‘yes’ ratio is the ratio of yes answers generated by the model.
@@
In this context, inner inconsistency refers to model responding ‘yes’ to even contradictory questions. An example is presented in Table 4. Despite the evident contradiction between the two questions (different $n$ setting), the model answer ‘yes’ to both. In fact, by comparing prompt I and prompt II, model shows this inner inconsistency at a ratio of 41%, which is very high because a model that randomly answers ‘yes’ or ‘no’ will bear an inner inconsistency of just about 25%. This high ratio reveals a significant challenge that it becomes uncertain what the model’s actual understanding is towards the question. Specifically, if the model answers ‘yes’ to a ground truth, we cannot confidently assert that the model precisely knows the number of objects in the given image. Moreover, even if the model answers ‘no’ to a hallucinated expression, it remains unclear whether the model accurately comprehends the ground truth or not. This inner inconsistency suggests that the model struggles to count the number of objects accurately and exhibits confusion, thereby contributing to number hallucination.
Input Picture Prompt and Output
[Uncaptioned image] Prompt: There are 2 motorcycles in this picture, is that correct? Answer yes or no. Answer: Yes
Prompt: There is 1 motorcycle in this picture, is that correct? Answer yes or no.
Answer: Yes
Table 4. Example of model answering ’yes’ to contradictory questions. This example is drawn from LLaVA.
@@ 4.2. Comparison Questions
Table 5. The results of comparison questions. Accuracy is calculated according to ground truth.
Input Picture Prompt and Output
[Uncaptioned image] Prompt: Which object has a larger number in this picture, backpack or dog? If they have the same number, answer same. Answer: Dog
Prompt: Which object has a larger number in this picture, dog or backpack? If they have the same number, answer same. Answer: Backpack
[Uncaptioned image] Prompt: Please select the correct option according to the given picture. A. There are more dogs than bicycles in this picture. B. There are more bicycles than dogs in this picture. C. The number of dogs and bicycles are the same in this picture. Answer: B
Prompt: Please select the correct option according to the given picture. A. There are more bicycles than dogs in this picture. B. There are more dogs than bicycles in this picture. C. The number of bicycles and dogs are the same in this picture. Answer: B
Table 6. Examples of model’s response changing after the order of objects (options) changes. The upper case shows one example of inner inconsistency of prompt I, while the lower case shows one example of inner inconsistency of prompt II. Both cases are drawn from MiniGPT-v2.
@@
Inner inconsistency is similar to what occurs in binary classification questions. In this context, inner inconsistency refers to the model providing different answers to the same question. By comparing within the same prompt format, it is surprising to find that after $c_{1}$ and $c_{2}$ are flipped, the model’s response sometimes changes. One example can be found in Table 6.
Our quantitative result shows that models generally output inconsistent answers for prompt I at a ratio of 25%, which is non-negligible. Prompt II also experiences this inconsistency even more severely (an example can be found in Table 6). For prompt II, the model responds with inconsistent answers of around 80%, which is disasterous. What’s more, if one only see the correct result(like the answers marked in green in) Table 6, he will probably fail to recognize the hidden hallucination inside the model, leading to oversight of the severity of hallucination.
This inconsistency reveals two facts. Firstly, the order of input and the format of prompt have a notable influence on the model’s output, making it challenging to evaluate the model’s performance with multiple-choice questions. Secondly, this inconsistency also reflects the model’s uncertainty about this question, which is a potential cause of number hallucination.
@@
Table 7. Results of finetuned models.
5. Mitigating Number Hallucination
5.1. Method
@@ 5.2. Experiment Setup
We utilize the MSCOCO2014 (Lin et al., 2014) training set as the foundational dataset and construct $S$ in the same manner as described in Section 3.2. We refrain from performing k-uniform sampling on the training set to avoid bias and retain as much data as possible. Since we have completely no idea about the data distribution in real-life scenarios, avoiding introducing prior biases and retaining as much data is more reasonable. This provides us with 233k pieces of data. For Direct method, all these data are used by directly asking like Section 3.1.
For Consistency(I), all these data are kept with prompt I, which provides us with 233k pieces of data similarly. The selection $m$ follows the same setting as setting III mentioned in Section 4.1. That is, 50% of training data sets $m=n_{c}$, while the rest of training data uniformly selects $m$ around $n$.
As for Consistency(II), 100k pieces of data are randomly selected with prompt II. This selection aims at limiting the total amount of data. By combining these two datasets, we get a dataset with 333k pieces of data, corresponding to Consistency(I+II). Note that all these consistency data are automatically constructed from the whole original MSCOCO2014 training set.
@@
We continue the experiments on LLaVA, InstructBLIP and MiniGPT-v2.444LLaVA and LLaVA-L bear the same architecture, so we only finetune LLaVA for simplicity. For simplicity and efficiency, we freeze the vision encoder as well as the LLM parts and only finetune the alignment part of each LVLM. For LLaVA, the alignment part is a two-layer MLP. For MiniGPT-v2, the alignment part is a linear layer. Regarding InstructBLIP, though the whole alignment part is a Q-former and a linear layer after it, we perform a preliminary experiment and observe that only finetuning the linear layer after Q-former instead of both Q-former and the linear layer offers a better performance on the Direct training method, so we also only finetune the linear layer after Q-former in InstructBLIP. Details of our hyper-parameters can be found in Appendix B.
5.3. Results and Analysis
The results are presented in Table 7. Some conclusions can be drawn from this result. Firstly, directly finetuning the model is useful for mitigating number hallucinations and provides a fine result.